Reports 1-1 of 1 Clear search Modify search
VIS (General)
satoru.ikeda - 13:19 Friday 20 August 2021 (17968) Print this report
Comment to Routine confirmation of DGS maintenance day (17831)

Routine confirmation of DGS maintenance day(Aug.20)

Performed routine check of DGS maintenance day.
No problems other than known problems.

Stepper motor
  Checked if the power supply can be controlled by BIO.
  Checked if Script can be started.
  Check that the limit switch installed is green.
 [Check result]
  Power control: No problem
  Limit SW: currently all NG, but will be adjusted in the future (known)

  Power control: No problem
  Limit SW: F0YAW is NG (known)

 Other than above: No problem

Release Sensor
 Confirmed that Release Sensor is 0. 1: Seated
 [Check result]
 ETMY, ITMX, ITMY : No problem
 ETMX_F1_1 : Abnormal, no other problems (known)

Images attached to this comment
Non-image files attached to this comment
Search Help